What are the three consecutive whole numbers whose sum is 441

Answer by macston(5194)

.
x, x+1, and x+2 are consecutive whole numbers
.
x+x+1+x+2=441
3x+3=441
3x=438
x=146
x+1=147
x+2=148
.
ANSWER: The three numbers are 146, 147, and 148
.
CHECK:
146+147+148=441
441=441